40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5066     Entry time: Sat Jul 30 05:11:45 2011
Author: Suresh 
Type: Update 
Category: Green Locking 
Subject: Y- end table clean-up 

The optics on the Y-end table which required to be moved have been repositioned.  Please see the attached pic for details.

The green beam is not yet aligned to the cavity. That is my next task.

Y-end_table_work.png

ELOG V3.1.3-